OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

Do you know what absolute value is?

OpenStudy (aum):

-3|x + 7| = -12 First step: Divide both sides by -3. What do you get?

OpenStudy (anonymous):

Remember, absolute value is the distance from 0

OpenStudy (anonymous):

dont forget at the end you will have two answers

OpenStudy (anonymous):

after i divide -3 on both sides i would get 4 as my answer then i would subtract 7 on both sides so my answer would be x= -3 then i would go and make 4 to -4 then subtract -4 on both sides and i would get x= -8

OpenStudy (anonymous):

well here -3|x + 7| = -12 -3lx + 7l/-3 = -12/-3 then you'll get two equations x + 7 = 4 x + 7 = -4 x = -3 x = -11